OpenStudy (anonymous):

Given the equation y - 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. Answer -3/4 x + y = 10 3x - 4y = -40 y = 3/4x + 12 y = 3/4 x + 10

OpenStudy (anonymous):

y - 4 = x + 8 y = x + 8 + 4 y = x + 12 this is slope intercept form (y = mx + b)